OpenStudy (anonymous):

x^2 = 1600tan^2(11)/[tan^2(27) + tan^2(11)] now how would you sqrt it what would x equal show the steps pls.

OpenStudy (anonymous):

@ganeshie8 @dan815 @ShadowLegendX @alekos @iambatman

OpenStudy (mimi_x3):

just sqrt \[x = sqrt(1600tan^2(11)....\] \[ x = \frac{40tan(11)}{tan(27)+tan(11)}\]

ganeshie8 (ganeshie8):

\[\large x^2 = \color{magenta}{\heartsuit} \] \[\large x = \pm \sqrt{ \color{magenta}{\heartsuit} }\]

OpenStudy (anonymous):

that cute heart :'(